Victor is y years old
His brother Fred is four years old than victor
The product of their age is 780
a) set up an equation to represent this information
b) solve your equation from a) to find victor's age

Answer by Boreal(15235)

y(y+4)=780
y^2+4y=780
y^2+4y-780=0
(y+30)(y-26)=0; Guess and check,start with 78/10, then 39/20, then look for two factors close together.
y=-30, 26, only the first is sensible.
Victor is 26 ANSWER
Fred is 30
The product is 780